# VideoLLaMA 3: Frontier Multimodal Foundation Models for Image and Video Understanding

### v3

@@ 3.2.3 Multi-task Fine-tuning
General LLaVA-SFT-665K (li2024llava, ), LLaVA-OV-SI (li2024llavaonevision, ), Cambrian-cleaned (tong2024cambrian, ), Pixmo (docs, cap, points, cap-qa, ask-model-anything) (molmo2024, ) 9.87M
Document DocVQA (mathew2021docvqadatasetvqadocument, ), Docmatix (laurençon2024building, ) 1.31M
Chart/Figure ChartQA (masry2022chartqa, ), MMC_Instruction (liu2023mmc, ), DVQA (kafle2018dvqa, ), LRV_Instruction (liu2023aligning, ), ChartGemma (masry2024chartgemmavisualinstructiontuningchart, ), InfoVQA (mathew2022infographicvqa, ), PlotQA (methani2020plotqa, ) 1.00M
OCR MultiUI (liu2024harnessingwebpageuistextrich, ), in-house data 0.83M
Grounding RefCoco (kazemzadeh2014referitgame, ), VCR (zellers2019vcr, ), in-house data 0.50M
Multi-Image Demon-Full (li2024fine, ), Contrastive_Caption (jiang2024mantisinterleavedmultiimageinstruction, ) 0.41M
Text-only Magpie (xu2024magpie, ), Magpie-Pro (xu2024magpie, ), Synthia (Synthia-70B-v1.2, ), Infinity-Instruct-subjective (InfinityInstruct2024, ), NuminaMath (li2024numinamath, ) 2.21M
@@
General LLaVA-Video-178K (zhang2024video, ), ShareGPT4o-Video (chen2024sharegpt4video, ), FineVideo (Farré2024FineVideo, ), CinePile (rawal2024cinepile, ), ShareGemini-k400 (sharegemini, ), ShareGemini-WebVID (sharegemini, ), VCG-Human (Maaz2024VideoGPT+, ), VCG-Plus (Maaz2024VideoGPT+, ), VideoLLaMA2 in-house data, Temporal Grounding in-house data 2.92M
@@
The general image data includes high-quality datasets, such as LLaVA-SFT-665K (liu2023improvedllava, ) and LLaVA-OV-SI (li2024llavaonevision, ), which serve as foundational resources for enhancing the model’s scene understanding. We also clean and filter the Cambrian-10M (tong2024cambrian, ) dataset. Furthermore, we incorporate meaningful data from the Pixmo dataset (molmo2024, ), including tasks such as document analysis, caption generation, and counting. These scene images cover a wide range of tasks, including captioning, counting, document understanding, mathematical reasoning, and etc.
@@
For OCR data, we consider two common cases in real-world scenarios: development scenarios and natural scenarios. For development scenarios, we use the MultiUI dataset (liu2024harnessingwebpageuistextrich, ) to activate the model’s capabilities in understanding and processing text within user interfaces. For natural scenarios, we leverage the Laion-OCR dataset to construct additional instruction-tuning data. The instruction-tuning data for OCR consists of the following five sub-tasks: 1) Text Existence Detection: Determine whether a specific piece of text exists within the image. 2) Text Localization: Locate a specific piece of text within the image and output its bounding box. 3) Text Recognition within a Bounding Box: Given a bounding box, recognize the text contained within it. 4) Text Comparison Between Images: Given two images, determine in which image the specified text appears. 5) Comprehensive Text Detection and Recognition: Detect and recognize all text present in the image.
For grounding images, we select data from established datasets such as RefCOCO (kazemzadeh2014referitgame, ) and VCR (zellers2019vcr, ), which focus on tasks of grounding visual elements in specific textual descriptions.
For multi-image scenes, we leverage the Demon-Full (li2024fine, ) and Contrastive-Caption (jiang2024mantisinterleavedmultiimageinstruction, ) datasets. The Demon-Full dataset is particularly valuable as it includes various tasks involving multi-image scenes, such as comparing differences between two images, generating captions for the final image in a comic strip, completing missing text in images with occluded portions, determining whether multiple images belong to the same category, and more. These tasks help the model handle complex scenarios involving multiple images, providing a more comprehensive understanding of how visual information can be interpreted across a series of related images. At the same time, such multi-image data further enhances the model’s video understanding capabilities.
For the video data used in this stage, we incorporate commonly used high-quality video caption datasets, along with a small amount of question-answering data. In addition, we supplement these with high-quality data from VideoLLaMA2 (damonlpsg2024videollama2, ) and in-house temporal grounding data. The in-house temporal grounding data specifically focuses on temporal relationships between video frames, enabling the model to grasp the sequence of events and understand the flow of actions across time. These combined data sources contribute to a more robust and nuanced video understanding capability for the model.
@@ 3.2.4 Video-centric Fine-tuning
  General Video LLaVA-Video-178K (zhang2024video, ), ShareGPT4o-Video (chen2024sharegpt4video, ), FineVideo (Farré2024FineVideo, ), CinePile (rawal2024cinepile, ), ShareGemini-k400 (sharegemini, ), ShareGemini-WebVID (sharegemini, ), VCG-Human (Maaz2024VideoGPT+, ), VCG-Plus (Maaz2024VideoGPT+, ), VideoRefer (yuan2024videorefer, ), VideoLLaMA2 in-house data, In-house synthetic data 3.03M
Streaming Video ActivityNet (krishna2017dense, ), YouCook2 (zhou2018towards, ), Ego4D-narration (grauman2022ego4d, ), Ego4D-livechat (chen2024videollm, ) 36.2K
@@
Temporal Grounding ActivityNet (krishna2017dense, ), YouCook2 (zhou2018towards, ), ViTT (huang2020multimodal, ), QuerYD (oncescu2021queryd, ), HiREST (zala2023hierarchical, ), Charades-STA (gao2017tall, ), Moment-10M (qian2024momentor, ), COIN (tang2019coin, ) 0.21M
Image-only LLaVA-SFT-665K (li2024llava, ), LLaVA-OV-SI (li2024llavaonevision, ) 0.88M
Text-only Magpie (xu2024magpie, ), Tulu 3 (lambert2024tulu3, ) 1.56M
@@
The video-centric fine-tuning stage is designed to tune VideoLLaMA3 to a video specialist and fully unleash its video understanding ability by focusing mainly on large-scale and high-quality video instruction following. We first collect videos with generally annotated caption, question, and answer from multiple open-source datasets including LLaVA-Video (zhang2024video, ), ShareGPT-4o (chen2024sharegpt4video, ), FineVideo (Farré2024FineVideo, ), CinePile (rawal2024cinepile, ), ShareGemini (sharegemini, ), VideoGPT+ (Maaz2024VideoGPT+, ) and VideoRefer yuan2024videorefer . These about 2.7M video-centric conversations eventually form a dataset across various scenes and tasks to serve as examples for teaching the model to understand complex dynamic and static content in videos.
@@
In addition, we further expand the data scale and strengthen the model by synthesizing dense captions and QAs of specific aspects. Specifically, following the pipeline proposed in (zhang2024video, ), we first filter 68K dynamic videos from Panda-70M (chen2024panda, ) dataset by optical flow, and then employ Qwen2-VL-72B (wang2024qwen2, ) to generate diverse dense captions and QAs for each video from the aspects of temporal understanding, spatial understanding, object description, and time-order understanding. Finally, 242K question-answer pairs are used for training.
@@
Besides general video-centric conversations, we also introduce the feature of streaming video understanding and temporal grounding to extend the application scenarios of our model. For streaming video understanding, we acquire data from ActivityNet (krishna2017dense, ), YouCook2 (zhou2018towards, ), and Ego4D (grauman2022ego4d, ), and organize video frames and multiple temporal dense captions in an interleaved manner as described in Section 3.1, aiming at enhancing the ability to understand fine-grained events in video and to sustain multi-turn conversations in streaming video. Since these videos are generally long, we cut them into small segments of up to two minutes according to the time interval of dense captions, and remove clips with overly dense and sparse captions. The synthetic streaming conversation from VideoLLM-Online (chen2024videollm, ) is also involved. For temporal grounding, we collect 205K data from datasets including ActivityNet (krishna2017dense, ), YouCook2 (zhou2018towards, ), ViTT (huang2020multimodal, ), QuerYD (oncescu2021queryd, ), HiREST (zala2023hierarchical, ), Charades-STA (gao2017tall, ), Moment-10M (qian2024momentor, ), and COIN (tang2019coin, ), and directly convert the grounding annotation to text format such as “1.0-2.0 s” for training.
@@
Finally, we emloy a certain amount of image-only and text-only data from LLaVA (li2024llava, ), LLaVA-OneVision (li2024llavaonevision, ), Magpie (xu2024magpie, ), and Tulu 3 (lambert2024tulu3, ) for mitigating the impact of catastrophic forgetting on the model’s capabilities.
@@ 3.3 Implementation Details
In this part, we briefly introduce the implementation details of each training stage. For all stages, we adopt the cosine learning rate scheduler. The warm up ratio of the learning rate is set as $0.03$. The maximum token length is set as $16384$, while the maximum token length for vision tokens is set as $10240$. In the stage of Vision Encoder Adaptation, when training VideoLLaMA3-2B, we initialize the vision encoder with the pre-trained weights of SigLIP (Zhai2023SigmoidLF, ) and the LLM with the pre-trained weights of Qwen2.5-2B yang2024qwen2 . For VideoLLaMA3-7B, the vision encoder is initialized with the fine-tuned SigLIP weights in VideoLLaMA3-2B and the LLM is initialized with Qwen2.5-7B yang2024qwen2 . The projector is implemented as a two-layer MLP with GELU as the activation function. In this stage, we only train the vision encoder and projector, and their learning rates are set as $1.0\times 10^{-5}$ and $1.0\times 10^{-3}$, respectively. For the remaining stages, the learning rates for the LLM, the projector, and the vision encoder are set as $1.0\times 10^{-5}$, $1.0\times 10^{-5}$, $2.0\times 10^{-6}$, respectively. The differential frame pruner is applied in the multi-task fine-tuning stage and the video-centric fine-tuning stage where video data is involved. The threshold to discard similar visual tokens is $0.1$. To limit context length, the visual tokens of videos are spatially downsampled after vision encoder by a factor of $2$ using bilinear interpolation. The visual tokens of images are only downsampled in the video-centric fine-tuning stage to align with video data. For loading video data, we first sample frames at $1$ frame per second using FFmpeg. These frames will be further sampled uniformly if the total number of frames is greater than a certain value, which is set to $180$ to accommodate most videos that last less than $3$ minutes.
@@ 4.1.1 Baselines
To comprehensively evaluate the image performance of VideoLLaMA3, we compare it against a diverse set of baselines. For the 2B version of the model, we select several strong methods, including SmolVLM (smolvlm2023, ), InternVL2.5-2B (internvl2.5, ), and Qwen2VL-2B (wang2024qwen2, ). For the 7B model, there are more options available. We choose to compare against Molmo-7B-D (molmo2024, ), InternVL2.5-8B (internvl2.5, ), LLaVA-OneVision (li2024llavaonevision, ), NVILA (liu2024nvila, ), and Qwen2VL-8B (wang2024qwen2, ).
@@ 4.1.2 Benchmarks
To evaluate VideoLLaMA3’s ability to understand various forms of texts in images, including documents, charts, and scene text, we conduct assessments on a range of benchmarks. Specifically, we use: 1) DocVQA (mathew2021docvqa, ) for document understanding, which evaluates the model’s ability to process and extract information from text in documents; 2) ChartQA (masry2022chartqa, ) and InfoVQA (mathew2021docvqa, ) for chart understanding, assessing the model’s ability to interpret and reason about data presented in graphical forms such as bar charts and line graphs; and 3) OCRBench (liu2023hidden, ) for scene text image understanding, which tests the model’s capacity to extract and comprehend text from images of real-world scenes.
@@
VideoLLaMA3’s mathematical reasoning capabilities are evaluated through the MathVista (lu2023mathvista, ) and MathVision (wang2024measuring, ) benchmarks. These benchmarks focus on evaluating the model’s ability to reason about and solve mathematical problems presented in visual formats, including text-based mathematical expressions and problem-solving tasks that require visual interpretation.
@@
To assess VideoLLaMA3’s ability to understand and reason about multiple images in conjunction, we evaluate the model on several widely used benchmarks, including MMMU-Pro (yue2024mmmupro, ), MMMU (yue2024mmmu, ), and BLINK (fu2025blink, ). These benchmarks test the model’s ability to draw connections between images, handle multiple visual inputs.
@@
Finally, to evaluate VideoLLaMA3’s performance in general question answering, particularly in real-world and complex scenarios, we conduct assessments using several challenging benchmarks. The benchmarks include: 1) RealWorldQA (realworldqa, ), which focuses on answering questions based on realistic images drawn from everyday scenarios, 2) AI2D (kembhavi2016diagram, ), which evaluates the model’s ability to reason about diagrams and science images, 3) GQA (hudson2019gqa, ), which assesses general question answering with a focus on complex visual reasoning tasks, and 4) MME (fu2023mme, ), which includes a wide variety of general knowledge questions that require a deep understanding of visual information.
@@ 4.1.3 Evaluation Protocols
When evaluating on benchmarks, we set the temparature as $0.0$. The maximum token length is set as the same as the training stage. For benchmarks involving the MCQ, we will give the prompt like “Answer with the option letter from the given choices directly.”. For the benchmarks with short answers, we will give the prompt like “Answer the question with a single word or phrase.”. We follow the original benchmarks to calculate the final scores, and we also align our evaluation protocols with other evaluation toolkits, such as lmms-eval zhang2024lmmsevalrealitycheckevaluation ; lmms_eval2024 and VLMEvalKit duan2024vlmevalkit .
@@ 4.2.1 Baselines
To comprehensively evaluate the video performance of VideoLLaMA3, we compare it with a diverse set of baseline models. Similar to image evaluation, there are few available models with a 2B parameter size in the community. We select several strong baselines, including Apollo-2B (zohar2024apollo, ), InternVL2.5-2B (internvl2.5, ), and Qwen2VL-2B (wang2024qwen2, ). For the 7B model, we compare it with generalist models such as Qwen2VL-7B (wang2024qwen2, ), InternVL2.5-8B (internvl2.5, ), and NVILA (liu2024nvila, ), as well as specialist models like LLaVA-Video (zhang2024video, ), Apollo-7B (zohar2024apollo, ), and our previous generation model, VideoLLaMA2 (damonlpsg2024videollama2, ).

### v4

@@ 3.2.3 Multi-task Fine-tuning
General LLaVA-SFT-665K [38], LLaVA-OV-SI [29], Cambrian-cleaned [39], Pixmo (docs, cap, points, cap-qa, ask-model-anything) [35] 9.87M
Document DocVQA [40], Docmatix [41] 1.31M
Chart/Figure ChartQA [42], MMC_Instruction [83], DVQA [84], LRV_Instruction [85], ChartGemma [86], InfoVQA [87], PlotQA [88] 1.00M
OCR MultiUI [89], in-house data 0.83M
Grounding RefCoco [90], VCR [91], in-house data 0.50M
Multi-Image Demon-Full [92], Contrastive_Caption [93] 0.41M
Text-only Magpie [94], Magpie-Pro [94], Synthia [95], Infinity-Instruct-subjective [82], NuminaMath [96] 2.21M
@@
General LLaVA-Video-178K [25], ShareGPT4o-Video [28], FineVideo [97], CinePile [98], ShareGemini-k400 [99], ShareGemini-WebVID [99], VCG-Human [22], VCG-Plus [22], VideoLLaMA2 in-house data, Temporal Grounding in-house data 2.92M
@@
The general image data includes high-quality datasets, such as LLaVA-SFT-665K [55] and LLaVA-OV-SI [29], which serve as foundational resources for enhancing the model’s scene understanding. We also clean and filter the Cambrian-10M [39] dataset. Furthermore, we incorporate meaningful data from the Pixmo dataset [35], including tasks such as document analysis, caption generation, and counting. These scene images cover a wide range of tasks, including captioning, counting, document understanding, mathematical reasoning, and etc.
@@
For OCR data, we consider two common cases in real-world scenarios: development scenarios and natural scenarios. For development scenarios, we use the MultiUI dataset [89] to activate the model’s capabilities in understanding and processing text within user interfaces. For natural scenarios, we leverage the Laion-OCR dataset to construct additional instruction-tuning data. The instruction-tuning data for OCR consists of the following five sub-tasks: 1) Text Existence Detection: Determine whether a specific piece of text exists within the image. 2) Text Localization: Locate a specific piece of text within the image and output its bounding box. 3) Text Recognition within a Bounding Box: Given a bounding box, recognize the text contained within it. 4) Text Comparison Between Images: Given two images, determine in which image the specified text appears. 5) Comprehensive Text Detection and Recognition: Detect and recognize all text present in the image.
For grounding images, we select data from established datasets such as RefCOCO [90] and VCR [91], which focus on tasks of grounding visual elements in specific textual descriptions.
For multi-image scenes, we leverage the Demon-Full [92] and Contrastive-Caption [93] datasets. The Demon-Full dataset is particularly valuable as it includes various tasks involving multi-image scenes, such as comparing differences between two images, generating captions for the final image in a comic strip, completing missing text in images with occluded portions, determining whether multiple images belong to the same category, and more. These tasks help the model handle complex scenarios involving multiple images, providing a more comprehensive understanding of how visual information can be interpreted across a series of related images. At the same time, such multi-image data further enhances the model’s video understanding capabilities.
For the video data used in this stage, we incorporate commonly used high-quality video caption datasets, along with a small amount of question-answering data. In addition, we supplement these with high-quality data from VideoLLaMA2 [46] and in-house temporal grounding data. The in-house temporal grounding data specifically focuses on temporal relationships between video frames, enabling the model to grasp the sequence of events and understand the flow of actions across time. These combined data sources contribute to a more robust and nuanced video understanding capability for the model.
@@ 3.2.4 Video-centric Fine-tuning
  General Video LLaVA-Video-178K [25], ShareGPT4o-Video [28], FineVideo [97], CinePile [98], ShareGemini-k400 [99], ShareGemini-WebVID [99], VCG-Human [22], VCG-Plus [22], VideoRefer [100], VideoLLaMA2 in-house data, In-house synthetic data 3.03M
Streaming Video ActivityNet [101], YouCook2 [102], Ego4D-narration [103], Ego4D-livechat [104] 36.2K
@@
Temporal Grounding ActivityNet [101], YouCook2 [102], ViTT [105], QuerYD [106], HiREST [107], Charades-STA [108], Moment-10M [109], COIN [110] 0.21M
Image-only LLaVA-SFT-665K [38], LLaVA-OV-SI [29] 0.88M
Text-only Magpie [94], Tulu 3 [111] 1.56M
@@
The video-centric fine-tuning stage is designed to tune VideoLLaMA3 to a video specialist and fully unleash its video understanding ability by focusing mainly on large-scale and high-quality video instruction following. We first collect videos with generally annotated caption, question, and answer from multiple open-source datasets including LLaVA-Video [25], ShareGPT-4o [28], FineVideo [97], CinePile [98], ShareGemini [99], VideoGPT+ [22] and VideoRefer [100]. These about 2.7M video-centric conversations eventually form a dataset across various scenes and tasks to serve as examples for teaching the model to understand complex dynamic and static content in videos.
@@
In addition, we further expand the data scale and strengthen the model by synthesizing dense captions and QAs of specific aspects. Specifically, following the pipeline proposed in [25], we first filter 68K dynamic videos from Panda-70M [112] dataset by optical flow, and then employ Qwen2-VL-72B [7] to generate diverse dense captions and QAs for each video from the aspects of temporal understanding, spatial understanding, object description, and time-order understanding. Finally, 242K question-answer pairs are used for training.
@@
Besides general video-centric conversations, we also introduce the feature of streaming video understanding and temporal grounding to extend the application scenarios of our model. For streaming video understanding, we acquire data from ActivityNet [101], YouCook2 [102], and Ego4D [103], and organize video frames and multiple temporal dense captions in an interleaved manner as described in Section 3.1, aiming at enhancing the ability to understand fine-grained events in video and to sustain multi-turn conversations in streaming video. Since these videos are generally long, we cut them into small segments of up to two minutes according to the time interval of dense captions, and remove clips with overly dense and sparse captions. The synthetic streaming conversation from VideoLLM-Online [104] is also involved. For temporal grounding, we collect 205K data from datasets including ActivityNet [101], YouCook2 [102], ViTT [105], QuerYD [106], HiREST [107], Charades-STA [108], Moment-10M [109], and COIN [110], and directly convert the grounding annotation to text format such as “1.0-2.0 s” for training.
@@
Finally, we emloy a certain amount of image-only and text-only data from LLaVA [38], LLaVA-OneVision [29], Magpie [94], and Tulu 3 [111] for mitigating the impact of catastrophic forgetting on the model’s capabilities.
@@ 3.3 Implementation Details
In this part, we briefly introduce the implementation details of each training stage. For all stages, we adopt the cosine learning rate scheduler. The warm up ratio of the learning rate is set as $0.03$. The maximum token length is set as $16384$, while the maximum token length for vision tokens is set as $10240$. In the stage of Vision Encoder Adaptation, when training VideoLLaMA3-2B, we initialize the vision encoder with the pre-trained weights of SigLIP [54] and the LLM with the pre-trained weights of Qwen2.5-2B [5]. For VideoLLaMA3-7B, the vision encoder is initialized with the fine-tuned SigLIP weights in VideoLLaMA3-2B and the LLM is initialized with Qwen2.5-7B [5]. The projector is implemented as a two-layer MLP with GELU as the activation function. In this stage, we only train the vision encoder and projector, and their learning rates are set as $1.0\times 10^{-5}$ and $1.0\times 10^{-3}$, respectively. For the remaining stages, the learning rates for the LLM, the projector, and the vision encoder are set as $1.0\times 10^{-5}$, $1.0\times 10^{-5}$, $2.0\times 10^{-6}$, respectively. The differential frame pruner is applied in the multi-task fine-tuning stage and the video-centric fine-tuning stage where video data is involved. The threshold to discard similar visual tokens is $0.1$. To limit context length, the visual tokens of videos are spatially downsampled after vision encoder by a factor of $2$ using bilinear interpolation. The visual tokens of images are only downsampled in the video-centric fine-tuning stage to align with video data. For loading video data, we first sample frames at $1$ frame per second using FFmpeg. These frames will be further sampled uniformly if the total number of frames is greater than a certain value, which is set to $180$ to accommodate most videos that last less than $3$ minutes.
@@ 4.1.1 Baselines
To comprehensively evaluate the image performance of VideoLLaMA3, we compare it against a diverse set of baselines. For the 2B version of the model, we select several strong methods, including SmolVLM [37], InternVL2.5-2B [34], and Qwen2VL-2B [7]. For the 7B model, there are more options available. We choose to compare against Molmo-7B-D [35], InternVL2.5-8B [34], LLaVA-OneVision [29], NVILA [36], and Qwen2VL-8B [7].
@@ 4.1.2 Benchmarks
To evaluate VideoLLaMA3’s ability to understand various forms of texts in images, including documents, charts, and scene text, we conduct assessments on a range of benchmarks. Specifically, we use: 1) DocVQA [113] for document understanding, which evaluates the model’s ability to process and extract information from text in documents; 2) ChartQA [42] and InfoVQA [113] for chart understanding, assessing the model’s ability to interpret and reason about data presented in graphical forms such as bar charts and line graphs; and 3) OCRBench [114] for scene text image understanding, which tests the model’s capacity to extract and comprehend text from images of real-world scenes.
@@
VideoLLaMA3’s mathematical reasoning capabilities are evaluated through the MathVista [115] and MathVision [116] benchmarks. These benchmarks focus on evaluating the model’s ability to reason about and solve mathematical problems presented in visual formats, including text-based mathematical expressions and problem-solving tasks that require visual interpretation.
@@
To assess VideoLLaMA3’s ability to understand and reason about multiple images in conjunction, we evaluate the model on several widely used benchmarks, including MMMU-Pro [117], MMMU [118], and BLINK [119]. These benchmarks test the model’s ability to draw connections between images, handle multiple visual inputs.
@@
Finally, to evaluate VideoLLaMA3’s performance in general question answering, particularly in real-world and complex scenarios, we conduct assessments using several challenging benchmarks. The benchmarks include: 1) RealWorldQA [120], which focuses on answering questions based on realistic images drawn from everyday scenarios, 2) AI2D [121], which evaluates the model’s ability to reason about diagrams and science images, 3) GQA [122], which assesses general question answering with a focus on complex visual reasoning tasks, and 4) MME [123], which includes a wide variety of general knowledge questions that require a deep understanding of visual information.
@@ 4.1.3 Evaluation Protocols
When evaluating on benchmarks, we set the temparature as $0.0$. The maximum token length is set as the same as the training stage. For benchmarks involving the MCQ, we will give the prompt like “Answer with the option letter from the given choices directly.”. For the benchmarks with short answers, we will give the prompt like “Answer the question with a single word or phrase.”. We follow the original benchmarks to calculate the final scores, and we also align our evaluation protocols with other evaluation toolkits, such as lmms-eval [124, 125] and VLMEvalKit [126].
@@ 4.2.1 Baselines
To comprehensively evaluate the video performance of VideoLLaMA3, we compare it with a diverse set of baseline models. Similar to image evaluation, there are few available models with a 2B parameter size in the community. We select several strong baselines, including Apollo-2B [15], InternVL2.5-2B [34], and Qwen2VL-2B [7]. For the 7B model, we compare it with generalist models such as Qwen2VL-7B [7], InternVL2.5-8B [34], and NVILA [36], as well as specialist models like LLaVA-Video [25], Apollo-7B [15], and our previous generation model, VideoLLaMA2 [46].